# Anomalous current fluctuations from Euler hydrodynamics

### v2

@@ Abstract
We consider the hydrodynamic origin of anomalous current fluctuations in a family of stochastic charged cellular automata. Using ballistic macroscopic fluctuation theory, we study both typical and large fluctuations of the charge current and reproduce microscopic results which are available for the deterministic single-file limit of the models. Our results indicate that in general initial fluctuations propagated by Euler equations fully characterize anomalous fluctuations on both scales. In the stochastic case, we find an additional contribution to typical fluctuations and conjecture the functional form of the typical probability distribution, which we confirm by numerical simulations.
Introduction.—Statistical physics aims to describe universal behaviors emerging from interactions of a large number of microscopic degrees of freedom. While traditionally the main focus has been equilibrium physics, in recent years it has been appreciated that out-of-equilibrium properties of the system can also display rich universal behaviors. A prominent tool that describes such physics is macroscopic fluctuations theory (MFT) [1], which captures large fluctuations of driven diffusive systems. The validity of MFT has been verified for a variety of systems [2, 3, 4, 5] and also on quantum simulators [6], but large fluctuations in these systems usually behave in a regular way with finite scaled cumulants of observables. This naturally raises the question of whether there exists systems that break such regularity and support new dynamical phenomena.
Recent numerical studies have found dynamical criticality and anomalous spin fluctuations in integrable spin chains [7, 8], which have also been probed experimentally using superconducting qubits [9]. While the mechanism responsible for anomalous spin fluctuations remains to be clarified, an exact calculation of fluctuations in a deterministic charged cellular automaton [10] provided a basic example of dynamical criticality in a deterministic many-body system. The procedure has been generalized to a wider class of single-file dynamics, which was explored in Refs. [11, 12, 13, 14, 15, 16]. Furthermore, anomalous fluctuations have also been observed in the easy-axis regime of an integrable spin chain [8, 17] and other systems including Dirac fluids [18] and stochastic processes [19], hinting at a more general setting supporting anomalous fluctuations.
@@
The model.—We consider a family of SCCA on a one-dimensional lattice of $L\in 2\mathbb{N}$ sites. The lattice configuration at fixed time $t\in\mathbb{Z}$ is given by strings ${\bf s}^{t}\equiv s^{t}_{L}s^{t}_{L-1}\ldots s^{t}_{1}$ of symbols $s^{t}_{x}\in\{\emptyset,-,+\}$, which correspond to vacancies and negatively or positively charged particles, respectively. Local symbol dynamics are given by a one-parameter stochastic two-body map $(s_{L}^{\prime},s_{R}^{\prime})=\Phi(s_{L},s_{R})$ which encodes the following dynamical rules (,)(,)($\emptyset$,$\emptyset$)$\rightarrow$($\emptyset$,$\emptyset$)( ∅ , ∅ ) → ( ∅ , ∅ ), $(\emptyset,c)\leftrightarrow(c,\emptyset)$ while $(c,c^{\prime})\to(c^{\prime},c)$ with probability $0\leq\Gamma\leq 1$ and $(c,c^{\prime})\to(c,c^{\prime})$ with probability $\overline{\Gamma}\equiv 1-\Gamma$ where $c,c^{\prime}\in\{-,+\}$, see Figure 1. We note that the dynamics can be understood as a composition of a stochastic six-vertex model in the particle-particle sector with free dynamics in the remaining sectors, see the Supplemental Material (SM) [23] for details.
The many-body dynamics is realized as a ‘brickwork’ circuit obtained by imposing periodic boundary conditions, $s^{t}_{x}=s^{t}_{L+x}$, and coupling alternating sites as ${\bf s}^{2t+2}=\Phi^{\rm odd}({\bf s}^{2t+1})$ and ${\bf s}^{2t+1}=\Phi^{\rm even}({\bf s}^{2t})$, where
@@
While the SCCA are super-integrable and support an exponential (in $L$) number of local conserved quantities [26], we presently restrict our study to a (closed) subset of conserved charges, comprising the number of right and left moving particles $\hat{Q}_{\pm}$ and the total charge $\hat{Q}_{c}$, all of which are given as sums of local densities $\hat{Q}_{i}({\bf s})=\sum_{x=1}^{L}\hat{q}^{i}_{x,t}({\bf s})$ where $i\in\mathcal{I}\equiv\{+,-,c\}$ and $\hat{\bullet}$ indicates a microscopic quantity, see SM [23] for details on charge densities and discrete space-time continuity relations.
@@
Hydrodynamics.—The hydrodynamics of the system turns out to have a simple structure with densities of left and right moving particles $\varrho_{\pm}/2$ and charge $\varrho_{c}$ forming a closed set of hydrodynamic equations. To study charge transport (in the chosen GGE) it is then sufficient to consider the hydrodynamics of only these three charges unlike in a standard integrable systems where all charges must be included.
The three-mode hydrodynamic equations of the SCCA are obtained by evaluating the flux jacobian
@@
where three modes are labeled by $\rho_{1}=\rho_{+}/2$, $\rho_{2}=\rho_{-}/2$, $\rho_{3}=\rho_{c}$ and the charge velocity $v$ is given by $v=p/\rho$ with $\rho=\rho_{1}+\rho_{2}$, $p=\rho_{1}-\rho_{2}$ the particle density and momentum respectively (see SM [23] for the derivation) and $j_{i}[\underline{\rho}]$ is the average current defined via the continuity equation $\partial_{t}\varrho_{i}+\partial_{x}j_{i}[\underline{\varrho}]=0$. Here, $\varrho_{i}$ is a space-time dependent density and the underline implies that the quantity is multicomponent. Charge inertness and chiral factorization of the dynamics amount to free ballistic propagation of the left and right movers $\partial_{t}(\varrho_{\pm}/2)\pm\partial_{x}(\varrho_{\pm}/2)=0$. On the other hand, charge transport is affected by particle dynamics and the Euler equation is of the form
$\partial_{t}\varrho_{c}+\partial_{x}(v\varrho_{c})=0.$ (3)
@@
A system of three-mode hydrodynamic equations similar to Eq. (3) along with the free chiral equations for ϱ±subscriptitalic-$\varrho_{\pm}$ was also recently derived for Dirac fluids by keeping only the most relevant terms based on symmetry considerations in Ref. [18]. In the present case, the hydrodynamic equations are obtained using the exact flux jacobian Eq. (2), and interestingly, are completely insensitive to the crossing probability $\Gamma$. That said, as we will see later, typical fluctuations do turn out to depend nontrivially on $\Gamma$. Before discussing this dependence, we first study large fluctuations of the SCCA, which can be fully characterized using only Euler hydrodynamic data.
@@
Anomalous charge FCS from BMFT.—The central object in the study of fluctuations is the generating function $\langle e^{{\rm i}\lambda\hat{J}(t)}\rangle=\int{\rm d}\lambda\,e^{{\rm i}\lambda J}\mathcal{P}(J|t)$ given by the probability distribution $\mathcal{P}(J|t)$ of time-integrated charge current $\hat{J}(t)=\int_{0}^{t}{\rm d}t^{\prime}\,\hat{\jmath}_{c}(0,t^{\prime})$. While the scale on which typical fluctuations take place $\hat{J}(t)\sim t^{1/{2z}}$ is set by the variance $\langle[\hat{J}(t)]^{2}\rangle^{c}\sim t^{1/z}$, large (atypical) fluctuations occur on the scale $\hat{J}(t)\sim t$ with a probability distribution governed by a large deviation principle $\mathcal{P}({J}=jt|t)\asymp e^{-tI(j)}$ where $I(j)$ is the large deviation rate function. The rate function is related to the scaled cumulant generating function (SCGF)
$F(\lambda)=\lim_{t\to\infty}t^{-1}\log\langle e^{\lambda\hat{J}(t)}\rangle$ (4)
via the Legendre transform $I(j)=\max_{\lambda\in\mathbb{R}}[\lambda j-F(\lambda)]$. We now compute the charge SCGF in the SCCA explicitly using BMFT.
The fundamental postulate of BMFT is that large fluctuations of the initial state are propagated according to Euler equations in space-time, thereby governing the large-scale dynamics at later times. The idea is implemented by writing the generating function as
@@
where the functional integral is performed over the space of functions supported on the space-time region $\mathbb{S}=\mathbb{R}\times[0,1]$, and the densities $\varrho_{i}(x,t)$ are the fluctuating mesoscopic densities evaluated at the Euler-scale coordinates $(\tau x,\tau t)$, i.e. $\varrho_{i}(x,t)=\hat{\varrho}_{i}(\tau x,\tau t)$. The auxiliary field $\underline{H}$ in Eq. (6) ensures that the fluctuating fields always satisfy the hydrodynamic equations, which in turn gives the BMFT action $S[\underline{\varrho},\underline{H}]$
@@
where $J(1)=\int_{0}^{1}{\rm d}t\,j_{c}(0,t)$ is the time-integrated current and $\mathcal{F}[\underline{\varrho}(\cdot,0)]$ is the rate function of initial fluctuating densities
with Lagrange multipliers $\beta^{i}(x,0)$, the free energy density $f(x,0)$ associated with the initial fluctuating density profile $\underline{\varrho}(\cdot,0)$, and the background free energy $f$. The saddle point of Eq. (5) now gives the BMFT equations
$\lambda\delta^{i}_{\leavevmode\nobreak\ c}\Theta(x)-\beta^{i}+\beta^{i}(x,0)-H^{i}(x,0)=0,$ (8)
$\lambda\delta^{i}_{\leavevmode\nobreak\ c}\Theta(x)-H^{i}(x,1)=0,$ (9)
$\partial_{t}\beta^{i}(x,t)+[\mathsf{A}^{\mathrm{T}}]^{i}_{\leavevmode\nobreak\ j}(x,t)\partial_{x}\beta^{j}(x,t)=0,$ (10)
$\partial_{t}H^{i}(x,t)+[\mathsf{A}^{\mathrm{T}}]^{i}_{\leavevmode\nobreak\ j}(x,t)\partial_{x}H^{j}(x,t)=0,$ (11)
where $\Theta(x)$ is the Heaviside step function. In terms of the solution $\underline{\beta}^{(\lambda)}(x,t)$ of the BMFT equations, the SCGF is computed as [21]
where $\underline{\varrho}^{(\lambda)}(x,t)=\underline{\varrho}[\underline{\beta}^{(\lambda)}(x,t)]$. Although solving BMFT equations is in general a formidable task even for (interacting) integrable systems, the simple hydrodynamic structure of the SCCA allows us to find their exact solution; see SM [23] for details. For $\beta^{\pm}=\beta$ and $t\in(0,1)$ the solutiong along the ray $x=0$ reads
$\beta^{(\lambda),\pm}(0,t)$ $=\beta-2\Theta(\mp\lambda b)\log\left[\frac{\cosh\beta^{c}(\lambda)}{\cosh\beta^{c}}\right],$ (14)
which, using Eq. (12), also gives the SCGF
$F(\lambda)=\frac{1}{2}\log\left[1+\Delta^{2}(\mu_{b}+\mu_{b}^{-1}-2)\right]$ (15)
where $\Delta^{2}=\rho(1-\rho)$ and $\mu_{b}=\cosh\lambda+|b|\sinh|\lambda|$. This coincides precisely with the microscopic result obtained in [10] up to prefactor $1/2$ due to a different convention for the unit of time.
As was also observed in Ref. [10], the SCGF Eq. (15) is anomalous in that the scaled higher cumulants $c_{2n>4}=\lim_{t\to\infty}t^{-1}\langle[\hat{J}(t)]^{2n}\rangle^{c}$ are singular. BMFT clarifies the physical origin of this anomalous behavior: since the normal mode $b(x,t)$ of the charge $\varrho_{c}(x,t)$ travels along the ray $x=0$, fluctuations along $x=0$ proliferate and thereby render clustering of connected charge current correlations algebraic in time [27]. While such a scenario generally results in superdiffusion [28], it is avoided here due to the charge conjugation symmetry of the system [11]. In general, we expect that, for a charge in the system to show anomalous fluctuations, it has to have a vanishing velocity in equilibrium (usually afforded by a $\mathbb{Z}_{2}$ symmetry of the charge) and other modes should be independent of their own velocities (i.e. no self-coupling), which would otherwise induce superdiffusion of the charge as in the heat mode with the $5/3$-Levy exponent in anharmonic chains [28]. It is also reasonable to assume that anomalous fluctuations are robust away from charge inertness, but a more precise analysis of such a situation is beyond the scope of the present Letter.
To our knowledge, the above solution is the first exact solution of full BMFT equations in an interacting model and therefore serves as an independent check of their validity. Furthermore, Eqs. (13) and (14) match the solution of the flow equation in ballistic fluctuation theory (see SM) [27, 29], corroborating the identification advocated in Ref. [21].
@@
Figure 2: Finite-time typical distributions of the integrated charge current $t^{1/4}\mathcal{P}^{[\Gamma]}(Jt^{-1/4}|t)$ (colored lines) in linear and logarithmic scales at different $\Gamma$ compared against distributions $\mathcal{P}^{[\Gamma]}_{\rm typ}$ (20) (black lines) with $\omega$ and $\sigma$ as fitting parameters, best-fit values reported in the SM [23]. Simulation parameters: $\rho=1/2$, $b=0$, $L=2^{20}$, $t_{\rm max}=2^{16}$, $5\times 10^{3}$ samples.
Typical fluctuations of deterministic dynamics from Euler hydrodynamics.—The probability distribution of typical charge fluctuations $\hat{J}(t)\sim t^{1/2z}$ is given by

### v3

@@ Abstract
We consider the hydrodynamic origin of anomalous current fluctuations in a family of stochastic charged cellular automata. Using ballistic macroscopic fluctuation theory, we study both large and typical fluctuations of the charge current and reproduce microscopic results available for the deterministic single-file limit of the models. Our results indicate that in general initial fluctuations propagated by Euler equations fully characterize anomalous fluctuations on both scales. In the stochastic case, we find an additional contribution to typical fluctuations and conjecture the functional form of the typical probability distribution, which agrees with numerical simulations.
I Introduction
Statistical physics aims to describe universal behaviors emerging from interactions of a large number of microscopic degrees of freedom. While traditionally the main focus has been equilibrium physics, in recent years it has been appreciated that out-of-equilibrium properties of the system can also display rich universal behaviors. A prominent tool that describes such physics is macroscopic fluctuations theory (MFT) [1], which captures large fluctuations of driven diffusive systems. The validity of MFT has been verified for a variety of systems [2, 3, 4, 5] and also on quantum simulators [6], but large fluctuations in these systems usually behave in a regular way with finite scaled cumulants of observables. This naturally raises the question of whether there exist systems that break such regularity and support new dynamical phenomena.
Recent numerical studies have found dynamical criticality and anomalous spin fluctuations in integrable spin chains [7, 8], which have also been probed experimentally using superconducting qubits [9]. While the mechanism responsible for anomalous spin fluctuations remains to be clarified, an exact calculation of fluctuations in a deterministic charged cellular automaton [10] provided a basic example of dynamical criticality in a deterministic many-body system. The procedure has been generalized to a wider class of single-file dynamics, which was explored in Refs. [11, 12, 13, 14, 15, 16]. Furthermore, anomalous fluctuations have also been observed in the easy-axis regime of an integrable spin chain [8, 17] and other systems including Dirac fluids [18] and stochastic processes [19], hinting at a more general setting supporting anomalous fluctuations. Anomalous spin transport in the isotropic Heisenberg chain has recently been under intense theoretical and experimental scrutiny [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30]. While the statistics of magnetization transfer is fully captured by the Baik-Rains subuniversality at the level of two-point quantities [31], higher-point quantities are distinct from the Kardar-Parisi-Zhang (KPZ) univerality class [32, 33, 34, 35] by symmetry [7], leading to partial KPZ physics.
@@ I Introduction
II The model
We consider a family of SCCA on a one-dimensional lattice of $L\in 2\mathbb{N}$ sites. The lattice configuration at fixed time $t\in\mathbb{Z}$ is given by strings ${\bf s}^{t}\equiv s^{t}_{L}s^{t}_{L-1}\ldots s^{t}_{1}$ of symbols $s^{t}_{x}\in\{\emptyset,-,+\}$, which correspond to vacancies and negatively or positively charged particles, respectively. Local symbol dynamics are given by a one-parameter stochastic two-body map $(s_{L}^{\prime},s_{R}^{\prime})=\Phi(s_{L},s_{R})$ which encodes the following dynamical rules (,)(,)($\emptyset$,$\emptyset$)$\rightarrow$($\emptyset$,$\emptyset$)( ∅ , ∅ ) → ( ∅ , ∅ ), $(\emptyset,c)\leftrightarrow(c,\emptyset)$ while $(c,c^{\prime})\to(c^{\prime},c)$ with probability $0\leq\Gamma\leq 1$ and $(c,c^{\prime})\to(c,c^{\prime})$ with probability $\overline{\Gamma}\equiv 1-\Gamma$ where $c,c^{\prime}\in\{-,+\}$, see Figure 1. The many-body dynamics is realized as a ‘brickwork’ circuit obtained by imposing periodic boundary conditions, $s^{t}_{x}=s^{t}_{L+x}$, and coupling alternating sites as ${\bf s}^{2t+2}=\Phi^{\rm odd}({\bf s}^{2t+1})$ and ${\bf s}^{2t+1}=\Phi^{\rm even}({\bf s}^{2t})$, where
@@ II The model
While the SCCA are super-integrable and support an exponential (in $L$) number of local conserved quantities [41], we presently restrict our study to a (closed) subset of conserved charges, comprising the number of right and left moving particles $\hat{Q}_{\pm}$ and the total charge $\hat{Q}_{c}$, all of which are given as sums of local densities $\hat{Q}_{i}({\bf s})=\sum_{x=1}^{L}\hat{q}^{i}_{x,t}({\bf s})$ where $i\in\mathcal{I}\equiv\{+,-,c\}$ and $\hat{\bullet}$ indicates a microscopic quantity. Explicitly, the local charge densities read $\hat{q}^{i}_{x,t}(\emptyset)=0$, $\hat{q}^{c}_{x,t}(\pm)=\pm 1$, $\hat{q}^{-}_{x,t}(\pm)=\delta_{x+t,{2\mathbb{Z}+1}}$ and $\hat{q}^{-}_{x,t}(\pm)=\delta_{x+t,{2\mathbb{Z}}}$ with corresponding local current densities $\hat{\jmath}^{i}_{x,t+1}=(-1)^{x}(\hat{q}^{i}_{x,t+1+(-1)^{x}}-\hat{q}^{i}_{x,t+1})/2$ that satisfy discrete continuity equations of the form
$\frac{1}{2}\left(\hat{q}^{i}_{x,t+2}-\hat{q}^{i}_{x,t}\right)+\hat{\jmath}^{i}_{x+1,t+1}-\hat{\jmath}^{i}_{x,t+1}=0.$ (2)
@@
III Hydrodynamics
The hydrodynamics of the system turns out to have a simple structure with densities of left and right moving particles $\varrho_{\pm}/2$ and charge $\varrho_{c}$ forming a closed set of hydrodynamic equations (here space-time dependent densities $\varrho_{i}$ are distinguished from the static ones $\rho_{i}$). To study charge transport (in the chosen GGE) it is then sufficient to consider the hydrodynamics of only these three charges unlike in standard integrable systems where all charges must be included.
The three-mode hydrodynamic equations of the SCCA are obtained by evaluating the flux Jacobian. To this end we first compute the static susceptibility matrix $\mathsf{C}_{ij}=L^{-1}\sum_{x,x^{\prime}=1}^{L}\langle\hat{q}^{i}_{x,t}\hat{q}^{j}_{x^{\prime},t}\rangle^{c}$, with $\langle xy\rangle^{c}\equiv\langle xy\rangle-\langle x\rangle\langle y\rangle$ as
$\mathsf{C}_{ij}=\frac{1}{2}\begin{bmatrix}\rho_{+}\overline{\rho}_{+}&0&b\rho_{+}\overline{\rho}_{+}\\ 0&\rho_{-}\overline{\rho}_{-}&b\rho_{-}\overline{\rho}_{-}\\ b\rho_{+}\overline{\rho}_{+}&b\rho_{-}\overline{\rho}_{-}&2\rho-2b^{2}(\rho^{2}+p^{2})\end{bmatrix},$ (3)
where $\rho\equiv(\rho_{+}+\rho_{-})/2$ and $p\equiv(\rho_{+}-\rho_{-})/2$ are the average particle and momentum density per lattice site. To compute the the static charge-current correlator $\mathsf{B}_{ij}=L^{-1}\sum_{x,x^{\prime}=1}^{L}\langle\hat{q}^{i}_{x,t}\,\hat{\jmath}^{j}_{x^{\prime},t+1}\rangle^{c}$ we observe that left/right movers move freely along diagonals of distinct sublattices with velocities ±$\pm 1$ which together with the symmetry $\mathsf{B}_{ij}=\mathsf{B}_{ji}$ immediately yields the relations $\mathsf{B}_{\pm\mp}=0,\ \mathsf{B}_{\pm\pm}=\pm\rho_{\pm}\overline{\rho}_{\pm}/2,\ \mathsf{B}_{\pm c}=\pm\rho_{\pm}\overline{\rho}_{\pm}b/2$. To evaluate the $\mathsf{B}_{cc}$ element, we note that the local dynamics generate a non-zero charge current only in the following six configurations (,±)(±,)plus-or-minusplus-or-minus($\emptyset$,$\pm$)$\leftrightarrow$($\pm$,$\emptyset$)( ∅ , ± ) ↔ ( ± , ∅ ), (+,)(,+)(+,-$\leftrightarrow$(-,+)( + , - ) ↔ ( - , + ). However, the latter two configurations enter into the average with identical ensemble weights while their current values are oppositely signed so that their contributions identically cancel. This also shows that $\mathsf{B}$ is independent of the crossing parameter $\Gamma$. Evaluating the remaining non-zero charge current configurations, we find $\mathsf{B}_{cc}=p(1-2b^{2}\rho)$, which gives the full static charge-current correlator
$\mathsf{B}_{ij}=\frac{1}{2}\begin{bmatrix}\rho_{+}\overline{\rho}_{+}&0&\rho_{+}\overline{\rho}_{+}b\\ 0&-\rho_{-}\overline{\rho}_{-}&-\rho_{-}\overline{\rho}_{-}b\\ \rho_{+}\overline{\rho}_{+}b&-\rho_{-}\overline{\rho}_{-}b&2p(1-2b^{2}\rho)\end{bmatrix}.$ (4)
With the matrices $\mathsf{B}$ and $\mathsf{C}$ at our disposal, we can obtain the flux Jacobian (5) by inverting the relation $\mathsf{B}=\mathsf{A}\mathsf{C}$, yielding
@@ III Hydrodynamics
where three modes are labeled by $\rho_{1}=\rho_{+}/2$, $\rho_{2}=\rho_{-}/2$, $\rho_{3}=\rho_{c}$ and the charge velocity $v$ is given by $v=p/\rho$ and $j_{i}[\underline{\rho}]$ is the average current defined via the continuity equation $\partial_{t}\varrho_{i}+\partial_{x}j_{i}[\underline{\varrho}]=0$. Here, $\varrho_{i}$ is a space-time dependent density and the underline implies that the quantity is multicomponent. Charge inertness and chiral factorization of dynamics amount to free ballistic propagation of the left and right movers $\partial_{t}(\varrho_{\pm}/2)\pm\partial_{x}(\varrho_{\pm}/2)=0$. On the other hand, charge transport is affected by particle dynamics and the Euler equation is of the form
$\partial_{t}\varrho_{c}+\partial_{x}j_{c}=0,\quad j_{c}=v\varrho_{c}.$ (6)
@@
A system of three-mode hydrodynamic equations similar to Eq. (6) along with the free chiral equations for ϱ±subscriptitalic-$\varrho_{\pm}$ was also recently derived for Dirac fluids by keeping only the most relevant terms based on symmetry considerations in Ref. [18]. In the present case, the hydrodynamic equations are obtained using the exact flux Jacobian Eq. (5), and interestingly, are completely insensitive to the crossing probability $\Gamma$. This being said, typical fluctuations appear to depend nontrivially on $\Gamma$. Before discussing this dependence, we first study large fluctuations of the SCCA, which can be fully characterized using only Euler hydrodynamic data.
IV Anomalous charge FCS from BMFT
@@
The central object in the study of fluctuations is the generating function $\langle e^{{\rm i}\lambda\hat{J}(t)}\rangle=\int{\rm d}\lambda\,e^{{\rm i}\lambda J}\mathcal{P}(J|t)$ given by the probability distribution $\mathcal{P}(J|t)$ of time-integrated charge current $\hat{J}(t)=\int_{0}^{t}{\rm d}t^{\prime}\,\hat{\jmath}_{c}(0,t^{\prime})$. While the scale on which typical fluctuations take place $\hat{J}(t)\sim t^{1/{2z}}$ is set by the variance $\langle[\hat{J}(t)]^{2}\rangle^{c}\sim t^{1/z}$, large (atypical) fluctuations occur on the scale $\hat{J}(t)\sim t$ with a probability distribution governed by a large deviation principle $\mathcal{P}({J}=jt|t)\asymp e^{-tI(j)}$ where $I(j)$ is the large deviation rate function. The rate function is related to the scaled cumulant generating function (SCGF) $F(\lambda)=\lim_{t\to\infty}t^{-1}\log\langle e^{\lambda\hat{J}(t)}\rangle$ via the Legendre transform $I(j)=\max_{\lambda\in\mathbb{R}}[\lambda j-F(\lambda)]$. We now compute the charge SCGF in the SCCA explicitly using BMFT.
BMFT is a hydrodynamic large deviation theory, which was introduced in Ref. [37] by generalizing the idea of MFT to systems that support ballistic transport, e.g. integrable systems. Following the principle of MFT, BMFT posits that large dynamical fluctuations at the Euler scale are entirely fixed by Euler hydrodynamics, namely fluctuating densities of the conserved charges transported by the Euler equation. However, the crucial difference from the standard MFT is that there is no source of fluctuations other than those in the initial condition at the Euler scale. This indicates that large dynamical fluctuations at later times can be obtained simply by propagating the initial fluctuations according to the Euler equation.
To implement this idea, we introduce the fluctuating mesoscopic densities evaluated at the Euler-scale coordinates $(\tau x,\tau t)$, i.e. $\varrho_{i}(x,t)=\hat{q}^{i}_{\tau x,\tau t}$ for $\tau\gg 1$. To determine how the fields $\varrho_{i}(x,t)$ fluctuate at later times, we need to characterize the initial fluctuations carried by $\varrho_{i}(x,0)$; the probability distribution of the initial fields is known to be given by $\mathbb{P}[\underline{\varrho}(\cdot,0)]\asymp e^{-\tau\mathcal{F}[\underline{\varrho}(\cdot,0)]}$, where $\mathcal{F}[\underline{\varrho}(\cdot,0)]$ is the density large deviation function (see Eq. (16) in Appendix. A). BMFT then asserts that the moment generating function $\langle e^{\lambda\hat{J}(\tau)}\rangle$ can be expressed as
@@
where the functional integral is performed over the space of functions supported on the space-time region $\mathbb{S}=\mathbb{R}\times[0,1]$. The auxiliary field $\underline{H}$ in Eq. (8) ensures that the fluctuating fields always satisfy the hydrodynamic equations, which in turn gives the BMFT action $S[\underline{\varrho},\underline{H}]$
@@
where $J(1)=\int_{0}^{1}{\rm d}t\,j_{c}(0,t)$ is the time-integrated current. The large time asymptotics of $\langle e^{\lambda\hat{J}(\tau)}\rangle$ is thus given by the saddle point of Eq. (7), and the solution to the BMFT equations that characterizes the saddle point, allows us to evaluate the SCGF exactly. In the SCCA it is possible to carry out the analysis explicitly (see Appendix. A for the derivation), resulting in the SCGF
$F(\lambda)=\tfrac{1}{2}\log\left[1+\Delta^{2}(\mu_{b}+\mu_{b}^{-1}-2)\right]$ (9)
where $\Delta^{2}=\rho(1-\rho)$ and $\mu_{b}=\cosh\lambda+|b|\sinh|\lambda|$. This coincides with the microscopic result obtained in [10] up to prefactor $1/2$ due to a different convention for the unit of time.
As also observed in Ref. [10], the generating function $\langle e^{{\rm i}\lambda\hat{J}(t)}\rangle$ is anomalous with singular higher scaled cumulants $c_{2n>4}=\lim_{t\to\infty}t^{-1}\langle[\hat{J}(t)]^{2n}\rangle^{c}$. BMFT clarifies the physical origin of this anomalous behavior: since the normal mode $b(x,t)$ of the charge $\varrho_{c}(x,t)$ travels along the ray $x=0$, fluctuations along $x=0$ proliferate and render clustering of connected charge current correlations algebraic in time [42]. While such a scenario generally results in superdiffusion [43], it is avoided here due to the charge conjugation symmetry of the system [11]. In general, we expect that, for a charge in the system to show anomalous fluctuations, it has to have a vanishing velocity in equilibrium (usually afforded by a $\mathbb{Z}_{2}$ symmetry of the charge) while other modes should be independent of their own velocities (i.e. no self-coupling), which would otherwise induce superdiffusion of the charge as in the heat mode with the $5/3$-Levy exponent in anharmonic chains [43].
@@
Figure 2: Finite-time typical distributions of the integrated charge current $t^{1/4}\mathcal{P}^{[\Gamma]}(Jt^{-1/4}|t)$ (colored lines) in linear and logarithmic scales at different $\Gamma$ compared against distributions $\mathcal{P}^{[\Gamma]}_{\rm typ}$ (14) (black lines) with $\omega$ and $\sigma$ as fitting parameters. Simulation parameters: $\rho=1/2$, $b=0$, $L=2^{20}$, $t_{\rm max}=2^{16}$, $5\times 10^{3}$ samples.
V Typical fluctuations of deterministic dynamics from Euler hydrodynamics
The probability distribution of typical charge fluctuations $\hat{J}(t)\sim t^{1/2z}$ is given by